# A Symmetric Multigrid-Preconditioned Krylov Subspace Solver for Stokes Equation

### v1

@@ 6.2.3 Brancher Example
To highlight the effectiveness of our preconditioner, we present a 3D brancher example. We keep the same boundary condition as 3D Poiseuille flow around cylinder example except that we substitute cylinder to multiple cuboids near the outflow boundary and use $\bar{u}=1.0m/s$. A 5-level multigrid is used within a $128^{3}$ unit cube. The sliced velocity field and the convergence results are shown in Fig. 14. We can see that our multiplicative preconditioner can converge very fast but the pure SQMR and multigrid method fails. This example shows the advantage of our approach in 3D case.
@@ 7 Conclusions
In this work, we propose and evaluate a novel multigrid preconditioner for solving the Stokes equations in both 2D and 3D scenarios. Our approach leverages the strengths of symmetric distributive smoother and Vanka smoother, which are well-suited for handling interior and boundary regions respectively. By combining these smoothers carefully, we achieve an efficient symmetric multigrid precondition which can be used by the SQMR method. Our numerical experiments in 2D and 3D have demonstrated the advantages of our approach. It not only gets the comparable convergence rates as pure multigrid methods, but also maintains fast convergence even when pure multigrid methods struggle with divergence. The application of our preconditioner as a robust and reliable solver, particularly in challenging 3D cases, shows its potential to significantly enhance the efficiency of solving the Stokes equations. Our future endeavors will focus on GPU acceleration to enhance the scalability of our methods, allowing for the efficient resolution of larger-scale problems.
@@ Appendix A Symmetry proof for symmetric distributive Gauss-Seidel
It’s obvious by using mathematical induction. ∎
@@
Substitute the first equation in Eq. 14 into the second one and denote $\mathbf{P}=\mathbf{M}\mathbf{\tilde{G}}^{-1}+\mathbf{\tilde{G}}^{-T}\mathbf{M}^{T}-\mathbf{\tilde{G}}^{-T}\mathbf{M}^{T}\mathbf{L}\mathbf{M}\mathbf{\tilde{G}}^{-1}$, we can get
@@
Appendix B Symmetry proof for addictive Vanka smoother
@@
We know that the series of addictive Vanka smoother can be written as
@@ Appendix C Symmetry proof for symmetric multiplicative Vanka smoother
Substitute the first equation in Eq. 20 into the second one and denote $\mathbf{P}=\mathbf{S}_{MV}^{(1)}+[\mathbf{S}_{MV}^{(1)}]^{T}-[\mathbf{S}_{MV}^{(1)}]^{T}\mathbf{L}\mathbf{S}_{MV}^{(1)}$, we can get
@@ Appendix D Symmetry proof for symmetric integration of smoother
For the second symmetric Vanka smoothing, $\mathbf{x}^{(0)}=[\mathbf{C}_{1}\mathbf{b}_{1},\mathbf{C}_{2}(\mathbf{b}_{2}-\mathbf{L}_{21}\mathbf{C}_{1}\mathbf{b}_{1})]^{T}$. By Lemma 2 and Eq. 22, we know that
@@ Acknowledgements
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author(s) used chatgpt in order to polish the sentences. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.

### v3

@@ 6.2.3 Brancher Example
To highlight the effectiveness of our preconditioner, we present a 3D brancher example. We keep the same boundary condition as 3D Poiseuille flow around cylinder example except that we substitute the cylinder with multiple cuboids near the outflow boundary and use $\bar{u}=1.0m/s$. A 5-level multigrid is used within a $128^{3}$ unit cube. The sliced velocity field and the convergence results are shown in Fig. 14. We can see that our multiplicative preconditioner converges very fast but the pure SQMR and multigrid method fails. This example shows the advantage of our approach in 3D case.
@@ 6.2.4 Porous Example
7 Stability and Efficiency Analysis
We present some additional experiments to assess the stability of our method under problems of different resolutions and conditioning, and multigrid cycle schemes.
7.1 Resolution and Conditioning
We seek to investigate the effect of domain resolution on the efficacy of our scheme. Although the increased spatial resolution does impact the condition number of the modified system matrix $\mathbf{\tilde{L}}$ used in multigrid, a healthy multi-resolution scheme would be expected to be consistently convergent, largely independent of resolution, even if the condition number itself increases. We illustrate the convergence of the 2D driven cavity example at different resolutions as shown in Fig. 16. We observe consistent convergence rates in this refinement analysis. We should indicate that the geometric complexity of a domain is another factor that can certainly affect convergence, especially in the presence of fine geometric features that are not well resolved or consistently represented across resolutions.
Figure 16: Convergence of 2D Driven Cavity Example at Different Resolutions
7.2 Cycle Schemes
We also design experiments using different cycle schemes. Specifically, we compare the convergence rates by using W-cycle and V-cycle on $1024^{2}$ Driven Cavity Example and $2200\times 410$ Poiseuille Flow around Cylinder Example as shown in Fig. 17 and 18. The results indicate that the W-cycle can indeed accelerate convergence rates (independent of runtime) in the cases where even pure multigrid methods converge. However, the increased time required for the W-cycle does not fully compensate for the improved convergence rates. Additionally, the deep-level W-cycle scheme fails to converge in our $1024^{2}$ Poiseuille flow around a hollow square example and brancher example. This suggests that the V-cycle is a more reliable scheme and potentially economical in terms of runtime.
Figure 17: Different Cycle Schemes on 2D $1024^{2}$ Driven Cavity Example
Figure 18: Different Cycle Schemes on 2D $2200\times 410$ Poiseuille Flow around A Cylinder Example
@@ 8 Conclusions
In this work, we propose and evaluate a novel multigrid preconditioner for solving the Stokes equations in both 2D and 3D scenarios. Our approach leverages the strengths of a symmetric distributive smoother and Vanka smoother, which are well-suited for handling interior and boundary regions respectively. By combining these smoothers carefully, we achieve an efficient symmetric multigrid preconditioner which can be used by the SQMR method. We demonstrate the efficacy of our scheme in numerical experiments in both 2D and 3D scenarios. Our method achieves comparable convergence rates to pure multigrid methods, and also maintains fast convergence even when pure multigrid methods struggle with divergence. The application of our preconditioner as a robust and reliable solver, particularly in challenging 3D cases, shows its potential to significantly enhance the efficiency of solving the Stokes equations. Our future endeavors will focus on GPU acceleration to enhance the scalability of our methods, allowing for the efficient resolution of larger-scale problems.
@@ Appendix A Symmetry proof for symmetric distributive Gauss-Seidel
This conclusion is obvious by using mathematical induction. ∎
@@
Substitute the first equation in Equation 14 into the second one and denote $\mathbf{P}=\mathbf{M}\mathbf{\tilde{G}}^{-1}+\mathbf{\tilde{G}}^{-T}\mathbf{M}^{T}-\mathbf{\tilde{G}}^{-T}\mathbf{M}^{T}\mathbf{L}\mathbf{M}\mathbf{\tilde{G}}^{-1}$, and we get
@@
Appendix B Symmetry proof for additive Vanka smoother
@@
We know that the series of additive Vanka smoother can be written as
@@ Appendix C Symmetry proof for symmetric multiplicative Vanka smoother
Substitute the first equation in Equation 20 into the second one and denote $\mathbf{P}=\mathbf{S}_{MV}^{(1)}+[\mathbf{S}_{MV}^{(1)}]^{T}-[\mathbf{S}_{MV}^{(1)}]^{T}\mathbf{L}\mathbf{S}_{MV}^{(1)}$, we can get
@@ Appendix D Symmetry proof for symmetric integration of smoother
For the second symmetric Vanka smoothing, $\mathbf{x}^{(0)}=[\mathbf{C}_{1}\mathbf{b}_{1},\mathbf{C}_{2}(\mathbf{b}_{2}-\mathbf{L}_{21}\mathbf{C}_{1}\mathbf{b}_{1})]^{T}$. By Lemma 2 and Equation 22, we know that